# CleanSurvival: Automated data preprocessing for time-to-event models using reinforcement learning

### v1

@@ 6 Conclusion
In this paper, we have introduced CleanSurvival, the first ever $Q$-learning-based framework that automates data preprocessing for survival analysis, addressing the often underserved challenge of automation in presence of censored data. We demonstrated its ability to enhance predictive performance and computational efficiency across different settings, compared to conventional methods or heuristic approaches.
@@
7 Acknowledgement
We are grateful for the input of Sergey Redyuk, whose suggestions greatly improved the manuscript. This work was supported by the German Federal Ministry of Education and Research via the project “Eventful” under grant ID 01IW23005.

### v2

@@ 6 Conclusion
In this paper, we have introduced CleanSurvival, a $Q$-learning-based framework that automates data preprocessing for survival analysis, addressing the often underserved challenge of automation in the presence of censored data. We demonstrated its ability to enhance predictive performance and computational efficiency across different settings, compared to conventional methods or heuristic approaches.
@@
7 Declarations
7.1 Ethics approval and consent to participate
Not applicable. This study used publicly available datasets.
7.2 Consent for publication
Not applicable.
7.3 Availability of data and materials
All datasets and code are available at https://github.com/phoenix0401/CleanSurvival.
7.4 Competing interests
The authors declare no competing interests.
7.5 Funding
This work was supported by the German Federal Ministry of Education and Research via the project “Eventful” under grant ID 01IW23005.
7.6 Authors’ contributions
YK conceptualized the package, wrote and implemented the code and wrote the manuscript. DS provided supervision, assisted in conceptualizing, writing and editing manuscript. GG helped in code optimization and implementation. SV provided supervision, funding acquisition and helped in reviewing and editing the manuscript. KP performed data evaluation, code implementation and editing the manuscript.
7.7 Acknowledgement
We are grateful for the input of Sergey Redyuk, whose suggestions greatly improved the manuscript.